# RNNs, CNNs and Transformers in Human Action Recognition: A Survey and A Hybrid Model

### v1

@@ 3.4 Discussion
In the field of action recognition, the choice of model – whether CNN-based, ViT-based, or a hybrid of CNN and ViT – significantly influences the outcome and efficiency of the task. CNN-based models are particularly adept at extracting local features due to their convolutional nature [120], making them highly effective in pattern recognition within images and videos. Their computational efficiency is a boon for real-time applications [121], and their robustness to input variations is notable [122]. However, CNNs often struggle with global contextual understanding [36] and are prone to overfitting. Moreover, their ability to model long-range temporal dependencies [123], which is crucial in action recognition, is somewhat limited.
@@
Hybrid models that combine CNNs and ViTs aim to harness the strengths of both architectures. They offer the local feature extraction capabilities of CNNs along with the global context awareness of ViTs, potentially providing a more balanced approach to action recognition. These models can be more efficient and versatile, adapting well to a range of tasks. However, this combination brings its own challenges, including increased architectural complexity, higher resource demands, and the need for careful tuning to balance the contributions of both CNN and ViT components. The choice among these models depends on the specific requirements of the action recognition task, such as the available computational resources, the nature and size of the dataset, and the types of actions that need to be recognized. For a summary of the advantages and disadvantages of these three architectural variations, see Table 4.
Table 4: Capability Comparison of Transformer-based, CNN-based, and Hybrid Models in Action Recognition
@@ 4 Proposed CNN-ViT Hybrid Architecture
In this section, we will detail our proposed CNN-ViT architecture for human action recognition, which leverages the benefits of both approaches described above. The architecture incorporates a TimeDistributed layer with a CNN backbone, followed by a ViT model to classify actions in video sequences.
Spatial Component (TimeDistributed CNN): The CNN backbone (i.e. Mobilenet [121]) in TimeDistributed layer processes each frame $\boldsymbol{X}\in\mathbb{R}^{K\times M\times N}$, and outputs spatial features vector $\boldsymbol{v}=p_{\theta}(\boldsymbol{X})\in\mathbb{R}^{L}$ where $p_{\theta}$ is the CNN model (e.g. MobileNet or VGG16) with parameters $\theta$ wrapped by TimeDistributed layer.
Let $\mathcal{X}$ be a collection of $n$ frames $\mathcal{X}=\{\boldsymbol{X}^{i}\}_{i=1}^{n}$, which is fed into the Spatial component of our architecture in Figure 7. The trained CNN $p_{\theta}$ is then used $n$ times to process $\mathcal{X}$ frame by frame and create a set of $n$ spatial feature vectors $\mathcal{V}=\{\boldsymbol{v}^{i}\}_{i=1}^{n}$, where $\boldsymbol{v}^{i}=p_{\theta}(\boldsymbol{X}^{i})$.
Temporal Component (ViT): In this proposed hybrid CNN-ViT model, the Vit is engineered to process sequences of $N$ vectors, see Figure 7, each representing a distinct frame of the video, but in the proposed ViT-only model, each vector is representing a distinct patch, see Figure 8. These vectors are first linearly projected into a high-dimensional space, facilitating the model’s ability to learn complex patterns within the data. To ensure the model captures the sequential nature of the input, positional encodings are added to these embeddings. The core of the ViT consists of two layers, each comprising a multi-head self-attention mechanism and a feed-forward network. The self-attention mechanism allows the model to weigh the importance of different patches relative to each other, while the feed-forward network, utilizing an Exponential Linear Unit (ELU) activation function, processes each position independently to capture global context. The ViT is designed to aggregate the information from all vectors and positional encodings into a single [CLS] token, which is prepended to the input sequence. The output vector associated with this [CLS] token, after propagation through the Transformer layers, serves as a comprehensive representation of the entire input, suitable for downstream classification tasks.
The sequence of vectors $\mathcal{V}$ which is outputted by the TimeDistributed CNN is then fed into the ViT. The ViT processes this sequence and outputs a final representation $\boldsymbol{z}$ which is represented as
$\boldsymbol{z}=\text{ViT}([\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{N}]).$ (16)
The final representation $\boldsymbol{z}$ is then fed into the Softmax layer to classify the action in the video. The Transformer encoder is designed to process a sequence of vectors, each representing one frame, and aggregate information into a single vector for classification.
@@
Figure 7: The hybrid CNN-ViT architecture.
@@
Figure 8: The ViT-only architecture.
5 Experiments
The goal of the presented experiments is not necessarily to produce a model that outperforms the state-of-the-art models in HAR field. Rather, the aim is to conduct a comparison among the ViT-only, CNN, and hybrid models to give further insights. The KTH dataset was chosen for its balanced representation of spatial and temporal features. Six experiments were conducted, training each of the aforementioned models on three different lengths of frame sequences. Care was taken to avoid pretraining in order to ensure the neutrality of the results. The Transnet model by Alomar et al. [126], was used to represent the CNN model. The ViT model was depicted in Figure 8. For the Spatial Component of the hybrid model, we employed the spatial component of Transnet and for the Temporal Component, we employed the same ViT model that we used in ViT-only model.
Experiment Settings. We constructed our model utilizing Python 3.6, incorporating the Keras deep learning framework, OpenCV for image processing, matplotlib, and the scikit-learn library. The training and testing were performed on a computer equipped with an Intel Core i7 processor, an NVidia RTX 2070 graphics card, and 64GB of RAM.
Data. In 2004, the Royal Institute of Technology unveiled the KTH dataset, a significant and publicly accessible dataset for action recognition [127]. Renowned as a benchmark dataset, it encompasses six types of actions: walking, jogging, running, boxing, hand-waving, and hand-clapping. The dataset features performances by twenty-five different individuals, introducing a diversity in execution. Additionally, the environment for each participant’s actions was deliberately altered, including settings such as outdoors, outdoors with scale changes, outdoors with clothing variations, and indoors. The KTH dataset comprises 2,391 video sequences, all recorded at 25 frames per second using a stationary camera against uniform backgrounds.
5.1 Results and Discussion
In the exploration of HAR within video sequences, three distinct models were evaluated i.e., CNN, ViT-only, and a Hybrid model. These models were trained on the KTH dataset, focusing on three different context lengths i.e., short (12 frames), medium (18 frames) and long (24 frames). The results from these experiments provide insightful revelations into the efficacy of each model under different temporal contexts, see Table 5.
Table 5: Experimental Results of three models on the KTH Dataset using three different context lengths. Hybrid-N refers to the model trained without pre-training whereas Hybrid-I is for the model pretrained on Imagenet[34].
Context length CNN-based ViT-only Hybrid-N Hybrid-I
@@ 5.1 Results and Discussion
In contrast, the ViT model demonstrated an improvement in performance with longer sequences, achieving accuracies of 92.44% for 12 frames, 92.82% for 18 frames, and 93.69% for 24 frames. This ascending pattern supports the notion that ViT architectures, with their inherent self-attention mechanisms, are well-suited to managing longer sequences. The ability of ViTs to assign varying degrees of importance to different parts of the sequence likely contributes to their enhanced performance on longer input frames.
The Hybrid CNN-ViT model showcased the highest and continuously improving accuracy rates across all frame lengths: 94.12% for 12 frames, 94.56% for 18 frames, and an impressive 95.78% for 24 frames. This model synergistically combines CNN’s robust spatial feature extraction capabilities with ViT’s efficient handling of temporal relationships via self-attention. The results from this model indicate that such a hybrid approach is particularly effective in capturing the complexities of action recognition tasks in video sequences, especially as the sequence length increases.
@@
Table 6: Comparison of proposed hybrid model with the state-of-the-art models on KTH dataset.
@@
Geng et al. [59] ICCSAE ’15 92.49
Arunnehru et al. [73] RoSMa ’18 94.90
Abdelbakya et al. [86] ITCE ’20 87.52
Jaouedi et al. [87] KSUCI journal ’20 96.30
Liu X et al [83] JAIHC ’20 91.93
Sahoo et al. [85] TETCI ’20 97.67
Lee et al. [91] CVF ’21 89.40
Basha et al. [96] MTA journal ’22 96.53
Ye et al. [118] CVF ’23 90.90
@@
Table 6 Shows that the presented CNN-ViT hybrid model achieved an impressive 97.89% accuracy on the KTH dataset, placing it prominently among state-of-the-art models for HAR. This performance is notably superior when compared to earlier benchmarks reported in the literature such as Geng et al. [59], who achieved 92.49%, and Arunnehru et al.[73] with 94.90%. Our model utilizes an ImageNet-pre-trained MobileNet [121] as the CNN backbone in the spatial component, which enhances its robust feature extraction capabilities. Combined with the dynamic attention mechanisms of ViT, it thereby enhances both the spatial and temporal processing of video sequences. Furthermore, the model not only surpasses other contemporary approaches like Liu X et al [83] (91.93%) and Lee et al. [91] (89.40%) but also shows competitive or superior performance against some of the highest accuracies in the field, such as Jaouedi et al. [87] (96.30%) and Basha et al. [96] (96.53%). Even in comparison to the high benchmark set by Sahoo et al. [85] (97.67%), our model demonstrates a marginal but significant improvement, underscoring the efficacy of integrating CNN with ViT. This integration not only facilitates more nuanced feature extraction across both spatial and sequential dimensions but also adapts more dynamically to the varied contexts inherent in video data, making it a potent solution for realistic action recognition scenarios.
The integration of CNN with ViT is particularly advantageous for enhancing feature extraction capabilities and focusing on relevant segments dynamically through the attention mechanisms of ViTs. This not only helps in improving accuracy but also in making the model more adaptable to varied video contexts, a key requirement for action recognition in realistic scenarios. This comparative advantage suggests that hybrid models are paving the way for future explorations in human action recognition, combining the best of convolutional and ViT-based architectures for improved performance and efficiency.
@@ 6 Challenges and Future Directions
Challenges The field of HAR faces several formidable challenges that stem from the inherent complexity of interpreting human movements within diverse and dynamic environments. One of the primary obstacles is the variability in human actions themselves, which can differ significantly in speed, scale, and execution from one individual to another [2]. This variability necessitates the development of sophisticated models capable of generalizing across a wide range of actions without sacrificing accuracy [128]. Additionally, the presence of complex backgrounds and environments further complicates the task of HAR. Systems must be adept at isolating and recognizing human actions against a backdrop of potentially distracting or obstructive elements, which can vary from the bustling activity of a city street to the unpredictable conditions of outdoor settings [129, 37].
Moreover, HAR systems must navigate the fine line between inter-class similarity and intra-class variability, where actions that are similar to each other (such as running versus jogging) require nuanced differentiation, while the same action can appear markedly different when performed by different individuals or under varying circumstances [130, 131]. The challenge of temporal segmentation adds another layer of complexity, as accurately determining the start and end of an action within a continuous video stream is crucial for effective recognition [132]. Coupled with the need for computational efficiency to process video data in real-time and the difficulties associated with obtaining large, accurately annotated datasets, these challenges underscore the multifaceted nature of HAR [133]. Addressing these issues is critical for advancing the field and enhancing the practical applicability of HAR systems in real-world applications, from surveillance and security to healthcare and entertainment.
Future Directions The motivation behind this work has been driven by the compelling need to bridge the existing gap between the spatial feature extraction capabilities inherent in CNNs and the dynamic temporal processing strengths found in ViTs [30]. Through the introduction of a novel hybrid model, an attempt has been made to leverage the synergistic potential of these technologies, thereby enhancing the accuracy and efficiency of HAR systems in capturing the complex spatial-temporal dynamics of human actions.
@@
Moreover, addressing the temporal segmentation challenge requires the integration of specialized temporal modules and sequence-to-sequence models to accurately determine the start and end of an action within continuous video streams. Computational efficiency remains paramount for real-time processing, necessitating ongoing efforts in model optimization and the exploration of synthetic data generation to mitigate the difficulties associated with obtaining large, accurately annotated datasets. Customizable Hybrid CNN-ViT models that can be tailored for specific applications, from surveillance to healthcare, will ensure that these advancements not only push the boundaries of academic research but also enhance practical applicability in real-world scenarios. Through these concerted efforts, Hybrid CNN-ViT models are poised to make significant contributions to the field of HAR, offering innovative solutions to its multifaceted challenges.
@@ 7 Conclusions
This survey provides a comprehensive overview of the current state of HAR by examining the roles and advancements of RNNs, CNNs, and ViTs. It delves into the evolution of these architectures, emphasizing their individual contributions to the field. The introduction of a hybrid model that combines the spatial processing capabilities of CNNs with the temporal understanding of ViTs represents a methodological advancement in HAR. This model aims to address the limitations of each architecture when used in isolation, proposing a unified approach that potentially enhances the accuracy and efficiency of action recognition tasks. The paper identifies key challenges and opportunities within HAR, such as the need for models that can effectively integrate spatial and temporal information from video data. The exploration of hybrid models, as suggested, offers a pathway for future research, particularly in improving model performance on complex video datasets. The discussion encourages further investigation into optimizing these hybrid architectures and exploring their applicability across various domains. This work sets a foundation for future studies to build upon, aiming to push the boundaries of what is currently achievable in HAR and to explore new applications of these technologies in real-world scenarios.

### v2

@@ 3.4 Discussion
In the field of HAR, the choice of models – whether CNN-based, ViT-based, or a hybrid of CNN and ViT – significantly influences the outcome and efficiency of the task. CNN-based models are particularly adept at extracting local features due to their convolutional nature [60], making them highly effective in pattern recognition within images and videos. Their computational efficiency is a boon for real-time applications [45], and their robustness to input variations is notable [82]. However, CNNs often struggle with global contextual understanding [88] and are prone to overfitting. Moreover, their ability to model long-range temporal dependencies [53], which is crucial in action recognition, is somewhat limited.
@@
Hybrid models that combine CNNs and ViTs aim to harness the strengths of both architectures. They offer the local feature extraction capabilities of CNNs along with the global context awareness of ViTs, potentially providing a more balanced approach to action recognition. These models can be more efficient and versatile, adapting well to a range of tasks. However, this combination brings its own challenges, including increased architectural complexity, higher resource demands, and the need for careful tuning to balance the contributions of both CNN and ViT components. The choice among these models depends on the specific requirements of the action recognition task, such as the available computational resources, the nature and size of the dataset, and the types of actions that need to be recognized.
For a summary of the advantages and disadvantages of these three architectural variations, see Table 6.
Table 6: Capability comparison between Transformer-based, CNN-based, and hybrid models in HARs.
@@ 4 Proposed CNN-ViT Hybrid Architecture
In this section, we present our proposed CNN-ViT architecture for HAR, leveraging the benefits of both approaches described in previous sections, see Figure 7. The architecture incorporates a TimeDistributed layer with a CNN backbone, followed by a ViT model to classify actions in video sequences.
Spatial component. Let $\mathcal{X}$ be a collection of $N$ frames, i.e., $\mathcal{X}=\{\boldsymbol{X}_{i}\}_{i=1}^{N}$. The CNN backbone (i.e. MobileNet in Howard et al. 45) in the TimeDistributed layer (see Figure 7) processes the indifivual frames $\boldsymbol{X}_{i}$ and outputs the spatial features vector $\boldsymbol{v}_{i}=p_{\theta}(\boldsymbol{X}_{i})\in\mathbb{R}^{L}$, where $p_{\theta}$ is the CNN model (e.g. MobileNet or VGG16) with parameters in $\theta$ wrapped by the TimeDistributed layer.
Temporal component. In the proposed hybrid CNN-ViT model, ViT is engineered to process the sequence of the $N$ spatial features vectors, i.e., $\{\boldsymbol{v}_{i}\}_{i=1}^{N}$, where each $\boldsymbol{v}_{i}$ represents a distinct frame of the input video clip, see Figure 7. Afterwards, the ViT block outputs a final representation $\boldsymbol{z}$, which is then fed into the softmax layer to classify the action in the video. In detail, the Transformer encoder is designed to process a sequence of vectors, each representing one frame, and aggregate information into a single vector for classification.
In the proposed ViT-only model in Figure 8 for the purpose of comparison, each vector represents a distinct patch. These vectors are first linearly projected into a high-dimensional space, facilitating the model’s ability to learn complex patterns within the data. To ensure the model captures the sequential nature of the input, positional encodings are added to these embeddings. The core of the ViT consists of two layers, each comprising a multi-head self-attention mechanism and a feed-forward network. The self-attention mechanism allows the model to weigh the importance of different patches relative to each other, while the feed-forward network, utilizing an exponential linear unit (ELU) activation function, processes each position independently to capture global context. The ViT is designed to aggregate the information from all vectors and positional encodings into a single [CLS] token, which is prepended to the input sequence. The output vector associated with this [CLS] token, after propagation through the Transformer layers, serves as a comprehensive representation of the entire input, suitable for downstream classification tasks.
@@
Figure 7: The hybrid CNN-ViT architecture for HARs.
@@
Figure 8: The ViT-only architecture for HARs.
4.1 Experiments
The goal of the presented experiments is not necessarily to produce a model that outperforms the state-of-the-art models in the HAR field. Rather, the aim is to conduct a comparison among the CNN, ViT-only, and hybrid models to give further insights.
The Royal Institute of Technology in 2004 unveiled the KTH dataset, a significant and publicly accessible dataset for action recognition [79]. The KTH dataset was chosen here for its balanced representation of spatial and temporal features. Renowned as a benchmark dataset, it encompasses six types of actions: walking, jogging, running, boxing, hand-waving, and hand-clapping. The dataset features performances by 25 different individuals, introducing a diversity in execution. Additionally, the environment for each participant’s actions was deliberately altered, including settings such as outdoors, outdoors with scale changes, outdoors with clothing variations, and indoors. The KTH dataset comprises 2,391 video sequences, all recorded at 25 frames per second using a stationary camera against uniform backgrounds.
Six experiments were conducted, with each of the aforementioned models trained on three different lengths of frame sequences. Care was taken to avoid pre-training in order to ensure the neutrality of the results. The TransNet model by [3] was adopted to represent the CNN model, and the ViT model was depicted in Figure 8. For the spatial component of the hybrid model, we employed the spatial component of TransNet; and for the temporal component, we employed the same ViT model that we used in the ViT-only model. We constructed our model utilizing Python 3.6, incorporating the Keras deep learning framework, OpenCV for image processing, matplotlib, and the scikit-learn library. The training and test were performed on a computer equipped with an Intel Core i7 processor, an NVidia RTX 2070 graphics card, and 64GB of RAM.
4.1.1 Results and Discussion
Table 7: Experimental results of different models on the KTH Dataset using three different context lengths. In particular, the hybrid model was trained without pre-training whereas Hybridpre is for the hybrid model pre-trained on ImageNet.
Context length CNN-based ViT-only Hybrid Hybridpre
@@ 4.1.1 Results and Discussion
Table 7 presents the quantitative results of the three distinct models, i.e., CNN, ViT-only, and a hybrid model on the KTH dataset, focusing on three different context lengths, i.e., short (12 frames), medium (18 frames), and long (24 frames). The results from these experiments provide insightful revelations into the efficacy of each model under different temporal contexts. More details are given below.
@@
In contrast, the ViT model demonstrated an improvement in performance with longer sequences, achieving accuracy of 92.44% for 12 frames, 92.82% for 18 frames, and 93.69% for 24 frames. This ascending pattern supports the notion that ViT architectures, with their inherent self-attention mechanisms, are well-suited to managing longer sequences. The ability of ViTs to assign varying degrees of importance to different parts of the sequence likely contributes to their enhanced performance on longer input frames.
The hybrid CNN-ViT model showcased the highest and continuously improving accuracy rates across all frame lengths: 94.12% for 12 frames, 94.56% for 18 frames, and an impressive 95.78% for 24 frames. Moreover, the pre-trained hybrid model showcased the same trend, with the best accuracy achieved. This type of model synergistically combines CNN’s robust spatial feature extraction capabilities with ViT’s efficient handling of temporal relationships via self-attention. The results from this model indicate that such a hybrid approach is particularly effective in capturing the complexities of action recognition tasks in video sequences, especially as the sequence length increases.
@@
Table 8: Comparison of the proposed hybrid model with the state-of-the-art models on the KTH dataset.
@@
[35] ICCSAE ’16 92.49
[5] RoSMa ’18 94.90
[1] ITCE ’20 87.52
[47] KSUCI journal ’20 96.30
[67] JAIHC ’20 91.93
[78] TETCI ’20 97.67
[62] CVF ’21 89.40
[8] MTA journal ’22 96.53
[109] CVF ’23 90.90
@@
To complete the comparison, Table 8 shows that the impressive 97.89% accuracy achieved by the presented CNN-ViT hybrid model on the KTH dataset places it prominently among state-of-the-art models for HAR. This performance is notably superior when compared to earlier benchmarks reported in the literature such as [35] with 92.49% and [5] with 94.90%. Our model utilizes an ImageNet-pre-trained MobileNet [45] as the CNN backbone in the spatial component, which enhances its robust feature extraction capabilities. Combined with the dynamic attention mechanisms of ViT, it can thereby enhance both the spatial and temporal processing of video sequences. Furthermore, our hybrid model not only surpasses other contemporary approaches like [67] (91.93%) and [62] (89.40%), but also shows competitive/superior performance against some of the highest accuracy in the field, such as [47] (96.30%) and [8] (96.53%). Even in comparison to the high benchmark set by [78] (97.67%), our hybrid model demonstrates a marginal but significant improvement, underscoring the efficacy of integrating CNN with ViT. This integration not only facilitates more nuanced feature extraction across both spatial and sequential dimensions but also adapts more dynamically to the varied contexts inherent in video data, making it a potent solution for realistic action recognition scenarios.
On the whole, the integration of CNN with ViT is particularly advantageous for enhancing feature extraction capabilities and focusing on relevant segments dynamically through the attention mechanisms of ViTs. This not only helps in improving accuracy but also in making the model more adaptable to varied video contexts, a key requirement for action recognition in realistic scenarios. This comparative advantage suggests that hybrid models are paving the way for future explorations in HAR, combining the best of convolutional and ViT-based architectures for improved performance and efficiency.
@@ 5 Challenges and Future Directions
The field of HAR faces several formidable challenges that stem from the inherent complexity of interpreting human movements within diverse and dynamic environments. One of the primary obstacles is the variability in human actions themselves, which can differ significantly in speed, scale, and execution from one individual to another [74]. This variability necessitates the development of sophisticated models capable of generalizing across a wide range of actions without sacrificing accuracy [73]. Additionally, the presence of complex backgrounds and environments further complicates the task of HAR. Systems must be adept at isolating and recognizing human actions against a backdrop of potentially distracting or obstructive elements, which can vary from the bustling activity of a city street to the unpredictable conditions of outdoor settings [97, 43].
HAR systems furthermore must navigate the fine line between inter-class similarity and intra-class variability, where actions that are similar to each other (such as running versus jogging) require nuanced differentiation, while the same action can appear markedly different when performed by different individuals or under varying circumstances [38, 113]. The challenge of temporal segmentation adds another layer of complexity, as accurately determining the start and end of an action within a continuous video stream is crucial for effective recognition [114]. Coupled with the need for computational efficiency to process video data in real-time and the difficulties associated with obtaining large, accurately annotated datasets, these challenges underscore the multifaceted nature of HAR [13]. Addressing these issues is critical for advancing the field and enhancing the practical applicability of HAR systems in real-world applications, from surveillance and security to healthcare and entertainment.
The motivation behind this work has been driven by the compelling need to bridge the existing gaps between the spatial feature extraction capabilities inherent in CNNs and the dynamic temporal processing strengths found in ViTs [4]. Through the introduction of a novel hybrid model, an attempt has been made to leverage the synergistic potential of these technologies, thereby enhancing the accuracy and efficiency of HAR systems in capturing the complex spatial-temporal dynamics of human actions.
@@
Moreover, addressing the temporal segmentation challenge requires the integration of specialized temporal modules and sequence-to-sequence models to accurately determine the start and end of an action within continuous video streams. Computational efficiency remains paramount for real-time processing, necessitating ongoing efforts in model optimization and the exploration of synthetic data generation to mitigate the difficulties associated with obtaining large and accurately annotated datasets. Customizable hybrid CNN-ViT models that can be tailored for specific applications, from surveillance to healthcare, will ensure that these advancements not only push the boundaries of academic research but also enhance practical applicability in real-world scenarios. Through these concerted efforts, hybrid CNN-ViT models are poised to make significant contributions to the field of HAR, offering innovative solutions to its multifaceted challenges.
@@ 6 Conclusions
This survey provides a comprehensive overview of the current state of HAR by examining the roles and advancements of CNNs, RNNs, and ViTs. It delves into the evolution of these architectures, emphasizing their individual contributions to the field. The introduction of a hybrid model that combines the spatial processing capabilities of CNNs with the temporal understanding of ViTs represents a methodological advancement in HAR. This model aims to address the limitations of each architecture when used in isolation, proposing a unified approach that potentially enhances the accuracy and efficiency of action recognition tasks. The paper identifies key challenges and opportunities within HAR, such as the need for models that can effectively integrate spatial and temporal information from video data. The exploration of hybrid models, as suggested, offers a pathway for future research, particularly in improving model performance on complex video datasets. The discussion encourages further investigation into optimizing these hybrid architectures and exploring their applicability across various domains. This work sets a foundation for future studies to build upon, aiming to push the boundaries of what is currently achievable in HAR and to explore new applications of these technologies in real-world scenarios.
[…]
Acknowledgements
K.A. and H.I.A. are thankful for the support from The Ministry of Education in Saudi Arabia and the Republic of Turkiye Ministry of National Education, respectively.
[…]
Author contributions Conceptualisation, K.A. and X.C.; methodology, K.A.; software, K.A.; validation,, all authors.; investigation, all authors; resources, K.A. and H.I.A.; data curation, K.A.; writing—original draft preparation, all authors; writing—review and editing, all authors; visualisation, K.A. and H.I.A.; supervision, X.C. All authors have read and agreed to the published version of the manuscript.
Declarations
Competing interests The authors declare no competing interests.